# Inference of entropy production for periodically driven systems

### v1

@@ V Analytical method to determine the intertransition time distribution
To calculate the intertransition time distribution $\psi_{\ell^{\prime}|\ell}(t)$ we define an auxiliary process that has absorbing states. This auxiliary process has an extra number of states that equals the number of transitions in $\mathcal{L}$, the set of visible transitions. All extra states are absorbing. For instance, if the trasition $\ell=1\to 2\in\mathcal{L}$ then the transition from $1$ to $2$ does not go to state $2$ in this new auxiliary process, it goes to a new absorbing state denoted $\ell=1\to 2$. The stochastic matrix associated with this auxiliary process is denoted $\mathbf{W}_{\text{aux}}(t)$.
@@
where the fourth state is associated with $1\to 2$ and the fifth state is associated with $2\to 1$.
@@
The conditional distribution $\mathcal{P}_{\ell^{\prime},\tau+t_{0}|\ell,t_{0}}$ is related to the time-dependent solution of this auxiliary process. Let’s denote a generic state states of the auxiliary process by $a$. The probability to be in state $a$ at time $t$ is denoted by $Q_{a}(t)$. The condition that $\ell$ and $t_{0}$ are given in the conditional probability $\mathcal{P}_{\ell^{\prime},\tau+t_{0}|\ell,t_{0}}$ is reflected in the initial probability of this auxiliary process. The time for the transition rates for this initial condition is $t_{0}$. If the transition $\ell=i\to j$, then the initial state of the auxiliary process is $j$, i.e., the initial probability distribution of the auxiliary process $Q^{\textrm{ini}}_{a}$ is a delta function that is one only for the state corresponding to $j$. Hence, we obtain $\mathcal{P}_{\ell^{\prime},\tau+t_{0}|\ell,t_{0}}=Q_{a}(\tau+t_{0})$, where the state $a$ is the absorbing corresponding to the transition $\ell^{\prime}$ and
$Q_{a}(\tau+t_{0})=\sum_{j}\left[\mathcal{T}\left\{\exp\int_{t_{0}}^{\tau+t_{0}}\mathrm{d}s\mathbf{W}_{\text{aux}}(s)\right\}\right]_{b,a}Q^{\textrm{ini}}_{b}(t_{0}).$ (30)
The intertransition time distribution $\psi_{\ell^{\prime}|\ell}(\tau)$ can be calculated by evaluating $Q_{a}(\tau+t_{0})$ from the above equation for all $t_{0}\in[0,T]$ and then applying equation (11). In Fig. 4, we show the agreement between the intertransition time distribution obtained analytically from this method and numerically from a single trajectory.
@@ VI Conclusion
We proposed a method for the inference of entropy production from the statistics of the intertransition times of a few visible transitions in periodically driven systems. We showed that it is possible to only look at the intertransition times by averaging out over all possible initial times. This procedure leads to an inequality between the real rate of entropy production $\sigma$ and our estimate $\hat{\sigma}$ that requires an extra term $X$. While our numerics indicate that an inequality $\sigma\geq\hat{\sigma}$ might be possible, the formal inequality we obtained here is $\sigma\geq\hat{\sigma}-X$. The emergence of this extra term and the necessity to generate two trajectories, one with forward protocol and another with backward protocol, are two main differences between the inference of entropy production for steady states and periodically driven systems.
@@
We applied the method to a simple three-state model for a molecular pump leading to two general lessons. First, even in a unicyclic system the estimate does not equal the exact rate of entropy production. This result shows a key difference in relation to steady states, for which the estimate does equal the exact entropy production in unicyclic systems van der Meer et al. (2022); Harunari et al. (2022). Second, the emergence of a net current in periodically driven systems, which is not a necessary condition for a non-zero entropy production, seems to be a necessary condition for $\hat{\sigma}$ to provide a meaningful estimate of the rate of entropy production. We observed that when the net current is zero the estimate is often numerically compatible with zero and does not capture any information about the real rate of entropy production. For cases with an emergent current, such as a molecular pump with a non-zero energy barrier or for a non-zero thermodynamic affinity, the method can provide a good estimate of the rate of entropy production. Hence, the method we propose here is a good candidate to estimate the rate of entropy production in physical systems with a non-zero net current such as molecular pumps and should not give much information for systems with a zero net current, such as heat engines.
@@
From a mathematical perspective it would be interesting to prove whether the inequality, observed to hold with our numerics, $\sigma\geq\hat{\sigma}$ indeed is correct. Interesting directions for future work include the study of inference of entropy production for the case of cyclic stochastic protocols, the investigation of other methods that do not require access to a trajectory with backward protocol, and a rigorous classification of physical systems for which the method can provide a good estimate.

### v2

@@ V Analytical method to determine the intertransition time distribution
To calculate the intertransition time distribution $\psi_{\ell^{\prime}|\ell}(t)$ we define an auxiliary process that has absorbing states. This auxiliary process has an extra number of states that equals the number of transitions in $\mathcal{L}$, the set of visible transitions. All extra states are absorbing. For instance, if the transition $\ell=1\to 2\in\mathcal{L}$ then the transition from $1$ to $2$ does not go to state $2$ in this new auxiliary process, it goes to a new absorbing state denoted $\ell=1\to 2$. The stochastic matrix associated with this auxiliary process is denoted $\mathbf{W}_{\text{aux}}(t)$.
@@
where the fourth state is associated with $2\to 1$ and the fifth state is associated with $1\to 2$.
@@
The conditional distribution $\mathcal{P}_{\ell^{\prime},\tau+t_{0}|\ell,t_{0}}$ is related to the time-dependent solution of this auxiliary process. Let’s denote a generic state of the auxiliary process by $a$. The probability of being in state $a$ at time $t$ is denoted by $Q_{a}(t)$. The condition that $\ell$ and $t_{0}$ are given in the conditional probability $\mathcal{P}_{\ell^{\prime},\tau+t_{0}|\ell,t_{0}}$ is reflected in the initial probability of this auxiliary process. The time for the transition rates for this initial condition is $t_{0}$. If the transition $\ell=i\to j$, then the initial state of the auxiliary process is $j$, i.e., the initial probability density of the auxiliary process $Q^{\textrm{ini}}_{a}(t_{0})$ is a delta function that is one only for the state corresponding to $j$ and zero otherwise. Hence, we obtain $\mathcal{P}_{\ell^{\prime},\tau+t_{0}|\ell,t_{0}}=\partial_{\tau}Q_{a}(\tau+t_{0})$, where the state $a$ is the absorbing state corresponding to the transition $\ell^{\prime}$ and
$Q_{a}(\tau+t_{0})=\sum_{b}\left[\mathcal{T}\left\{\exp\int_{t_{0}}^{\tau+t_{0}}\mathrm{d}s\mathbf{W}_{\text{aux}}(s)\right\}\right]_{b,a}Q^{\textrm{ini}}_{b}(t_{0}).$ (32)
The intertransition time distribution $\psi_{\ell^{\prime}|\ell}(\tau)$ can be calculated by evaluating $Q_{a}(\tau+t_{0})$ from the above equation for all $t_{0}\in[0,T]$, taking its derivative and then applying equation (11). In Fig. 6, we show the agreement between the intertransition time distribution obtained analytically from this method and numerically from a single trajectory.
@@ VI Conclusion
We proposed a method for the inference of entropy production from the statistics of the intertransition times of a few visible transitions in periodically driven systems. If sufficient data is available, the estimator $\sigma_{\mathcal{L}}$, which depends on both the intertransition time and the initial time of the first transition estimator, provides the tightest bound to the rate of entropy production. The more accessible estimator $\hat{\sigma}$ is obtained by eliminating the need to track the initial time, it is related to the rate of entropy production $\sigma$ with a formal inequality that requires the extra term $X$. While our numerics indicate the possibility of the inequality $\sigma\geq\hat{\sigma}$, this result remains restricted to our case study. The formal inequality we obtained here is $\sigma\geq\hat{\sigma}-X$. The emergence of this extra term and the necessity to generate two trajectories, one with forward protocol and another with backward protocol, are two main differences between the inference of entropy production for steady states and periodically driven systems.
@@
We applied our estimators to a simple three-state model for a molecular pump leading to two general lessons. First, even in a unicyclic system the estimate does not equal the exact rate of entropy production. This result shows a key difference in relation to steady states, for which the estimate does equal the exact entropy production in unicyclic systems van der Meer et al. (2022); Harunari et al. (2022). Second, the emergence of a net current in periodically driven systems, which is not a necessary condition for a non-zero entropy production, seems to be a necessary condition for $\hat{\sigma}$ to provide a meaningful estimate of the rate of entropy production. We observed that when the net current is zero the estimate is often numerically compatible with zero and does not capture any information about the real rate of entropy production. It would be interesting to investigate in which regimes the estimator $\hat{\sigma}$ is exactly zero, for instance in the case of no-pumping. For cases with an emergent current, such as a molecular pump with a non-zero energy barrier or for a non-zero thermodynamic affinity, the method can provide a good estimate of the rate of entropy production. Hence, the method we propose here is a good candidate to estimate the rate of entropy production in physical systems with a non-zero net current such as molecular pumps and should not give much information for systems with a zero net current, such as heat engines.
A recent independent study Maier et al. (2024) on the same topic of inference of entropy production in periodically driven systems came to our knowledge after we completed this work. Their approach is complementary to ours: They propose an estimator based on intertransition time statistics similar to $\hat{\sigma}$, in the form of a numerical conjecture, that has the advantage of only depending on statistics associated with the forward trajectory. Our work shows that, at the expense of running trajectories with the reverse protocol, a formal inequality between estimator, extra term $X$ and real rate of entropy production can be proven. Moreover, it would be interesting to compare the performance of our present bounds with the bound for periodically driven systems associated with the thermodynamic uncertainty relation from Koyuk et al. (2020), which depends on a single current. Such a comparison can be meaningful if the visible transitions are the ones related to a single current.
@@
From a mathematical perspective, it would be interesting to prove whether the inequality $\sigma\geq\hat{\sigma}$, observed to hold within our numerics, is indeed correct. Interesting directions for future work include the study of inference of entropy production for the case of cyclic stochastic protocols, the investigation of other methods that do not require access to a trajectory with backward protocol, and a rigorous classification of physical systems for which the method can provide a good estimate.
VII Acknowledgments
PH was supported by the project INTER/FNRS/20/15074473 funded by F.R.S.-FNRS (Belgium) and FNR (Luxembourg). CEF acknowledges the financial support from brazilian agencies CNPq and FAPESP under grants 2024/03763-0 and 2022/15453-0.